# Hybrid User Interfaces: Past, Present, and Future of Complementary Cross-Device Interaction in Mixed Reality

### v1

@@ 2.3 Identifying Hybrid User Interfaces
Research on HUIs shares several common themes (e.g., complementarity, heterogeneity), yet the term remains technology-driven and potentially misunderstood: A large number of prior HUIs describe the term as a combination of 2D and 3D technology [13, 38, 62], including the initial system demonstration by Feiner and Shamash [47]. Moreover, the term HUI was previously also used for different constellations of interaction components (e.g., desktop combined with tangibles [138], mobile devices [10], or conversational interfaces [93]), indicating wider applicability. For the purpose of our survey, we modernize the definition of HUIs by Feiner and Shamash [47] within the current research landscape:
@@
Hybrid User Interfaces are an area of cross-device computing that leverages distinct benefits of heterogeneous interaction components.
@@
In contrast to its initial definition, we explicitly concentrate on conceptual device capabilities instead of specific technologies to avoid being limited by current hardware capabilities, but are intentionally vague about possible device combinations (e.g., see cross-device interaction [16] for a potential ontology). To this end, we use the term interaction component throughout this work to refer to a standalone device set with the input and output capabilities necessary for interacting with a given application (e.g., desktop with mouse and keyboard, VR HWD with controllers).
@@
While this definition highlights the broad potential design space of HUIs, it also lacks specificity and, therefore, limits its utility for our survey: For example, a combination of tangibles and a desktop computer might be considered a HUI and likewise, a combination of an AR HWD with a tablet—increasing the blurriness of the research stream instead of increasing its focus. Therefore, to explore this vast design space and gain concrete insights into commonalities of HUIs, we focus our work on the specific HUI subset combining 2D with MR-enabled interaction components. This combination captures our interest, as it was initially presented by Feiner and Shamash [47], is most prevalent in prior HUI literature [125], and also reflects on work presented at the IEEE ISMAR workshop on HUIs (2023) [68], constituting the most recent and comprehensive outlet for work on HUIs. We thereby explicitly exclude the vast research area of tangible interaction [74] or nascent research areas (e.g., brain–computer interfaces [28], Internet of Things [122]) from this subset, thereby increasing the specificity of our survey.
@@ 2.4 Attributes of Mixed Reality Hybrid User Interfaces
We framed our work on the contemporary and most prevalent subset of HUIs—a combination of 2D interaction components with MR environments. We derive three attributes based on previous usage of the term to further guide our literature review. Please note that the first two attributes are indicative for the broader area of HUIs, while the last attribute specifies the constraints of MR-HUIs.
Multiple interaction components in heterogeneous roles.  HUIs typically combine multiple interaction components where each fulfills a need that is not adequately addressed by other components. Although each interaction component is self-contained, a HUI is deliberately spread across multiple components to intentionally “take advantage of the strong points of each” [47]. We also see the potential beyond technological aspects and broaden our scope to include combinations of heterogeneous roles.
Interaction components are codependent.  A HUI is composed of multiple interaction components, but the real power does not emerge from any individual interaction component, but from the interaction of all of them [149]. A HUI thus acts as one holistic application from the user’s perspective. The deliberate spread of responsibilities across interaction components requires a certain degree of co-dependency between interaction components (i.e., a system may become nonfunctional without all interaction components present). We examine both synchronous (i.e., using interaction components in parallel) and asynchronous usage (i.e., using interaction components in sequence) [13, 69].
Complementing 2D with mixed reality interaction components.  A common aspect found within prior HUI literature is to “pair 3D perception and direct 3D interaction with 2D system control and precise 2D interaction” [13]. Given the (at times) conflicting needs of 2D and 3D spaces, combining 2D and MR interaction components can yield a superior result. For the remainder of this work, we refer to such interaction components as either 2D interaction component or MR interaction component.
@@ 3 Review Methodology
We look at research that matches our previously specified attributes to obtain corpus of MR-HUI publications. This section describes the process for identifying, filtering, and analyzing relevant publications, following the PRISMA [108] guidelines. We present the general search strategy (Sec. 3.1), selection process (Sec. 3.2), and data extraction (Sec. 3.3). Moreover, the limitations of our procedure (Sec. 3.4) and possible extensions of the survey (Sec. 3.5) are highlighted. Finally, we describe how this survey can be used by other researchers and practitioners (Sec. 3.6). The complete survey corpus, its coding, and the scripts used to prepare and analyze the data can be found in the supplementary material.
Figure 2: Overview of our reviewing process including paper counts, following PRISMA. Records were rejected once a single exclusion criterium was fulfilled, yet, some records potentially fulfilled multiple criteria (e.g., survey papers [4]). The initial corpus of papers (n=579) including all codes is part of the supplemental material of this submission.
@@ 3.1 Search Strategy
Identification via databases.  Our survey focuses on publications that combine 2D interaction components with MR environments. As the term “HUI” is not consistently used in previous work, we used different adjacent terms and their synonyms (see Sec. 2.2) to build a query that captures a broader range of potential publications that present HUI applications.
@@
The query is composed of these three sets and was constructed as follows:
    Set A or (Set B and Set C)
We focused on archival, peer-reviewed publications, including full and short papers in journals and conferences, as well as book chapters, workshop submissions, posters, and works in progress. We searched in common digital libraries for publications in Human–Computer Interaction and Visualization, namely IEEE Xplore, ACM Digital Library, Springer Link, and Wiley Online Library.The searches with the described query [^4] The specific search syntax for each library can be found in the supplemental material. in these libraries [^5] At the time of performing this query, the Springer library did not support an advanced search. Thus, we filtered the results using a script (see supplemental material). resulted in 355 publications (see Fig. 2). The cut-off date for all searches was August 12th, 2024.
Identification via other methods.  In addition to the query search, we selected papers through three other methods (see Fig. 2). First, we selected all papers referring to the original definition of HUI by Feiner and Shamash [47] through a Google Scholar search using “Publish or Perish” [61]. Second, we manually added further publications that we found relevant. Finally, we selected all publications presented at the IEEE ISMAR 2023 Workshop on Hybrid User Interfaces [68]. The three methods resulted in 224 publications (see Fig. 2). The cut-off date was August 12th, 2024.
@@ 3.2 Selection Process
After retrieving the initial corpus of papers (n = 579), we preprocessed the different output formats and merged them in one table. Three authors screened the reviewed corpus and filtered out duplicates and papers unrelated to MR. In addition, workshop proposals, complete books, dissertations, or conference proceedings were also removed.
Subsequently, three authors checked the eligibility of the remaining publications (n = 312). For that, the following exclusion criteria (based on Sec. 2.4) were defined:
  • The publication is a duplicate (content-wise) of another (e.g., demonstration [113] of system [114]).
  • No prototype or system was presented (e.g., discussion of opportunities and challenges [157, 155, 86, 115]).
  • Only a technological basis is described (e.g., frameworks [65, 122]).
  • The publication only presents a summary of systems already in our corpus (e.g., position papers [45, 44]).
  • A pure MR system without other interaction components is described (e.g., combining AR and VR environments [26, 6]).
  • Other device capabilities are only used as input or output modality (i.e., interaction components are not standalone, e.g., attaching input sensors to an AR HWD [97, 31]).
  • The publication only presents a collaboration across asymmetric devices (e.g., one user on a 2D interaction component, another in an MR environment [105, 94]).
During both selection steps, we split the data set between authors and discussed entries if the decision was unclear. This resulted in 83 relevant publications within our corpus.
3.3 Data Extraction and Code Book
We first created an initial code book to extract data from the remaining 83 papers in our corpus. Since we see HUIs as a sub-category of cross-device interaction, we started with aspects defined by Brudy et al. [16] in their survey. This was extended with common HCI metadata, such as the contribution type defined by Wobbrock and Kientz [152] or the evaluation strategy outlined by Ledo et al. [88]. Furthermore, we recorded each publication’s challenges, future work, use case, and devices and terminology used. Once we defined an initial set of codes, we selected 10 papers at random from our literature corpus that three authors reviewed. This allowed us to (1) verify if the categorization works, (2) add missing codes, and (3) establish a common ground between authors. With the categories finalized, four authors coded the remaining papers. Two authors independently coded every publication. No author assessed the relevance of their own work. After each paper was coded, the same four authors discussed each publication and combined the two coding entries (e.g., clarifying conflicts). They further clustered challenges, use cases, and terminology used throughout the corpus.
@@ 3.4 Limitations and Further Considerations
Strict Eligibility Criteria.  We searched for a specific type of device combination (i.e., MR-enabling devices and conventional 2D displays), leading to strict eligibility criteria (see Sec. 2.4). To avoid being overly restrictive, we decided to integrate papers that meet our characteristics but leave room for interpretation, especially papers representing potential directions for future research. We discuss these edge cases and their implications in Sec. 4.3. We acknowledge that our corpus is not exhaustive with regard to possible HUI systems caused by the use of said criteria and our intentional focus on MR-HUI. Yet, we see our corpus as a representative and substantial set of research on MR-HUI, showcasing trends that can be generalized to the larger set of papers not captured by our query.
@@ 3.5 Dissemination and Extension
To allow others to benefit from our survey, we have made our literature corpus available using the Indy Survey Tool [29] as a GitHub project that hosts an interactive website for users to explore and filter our survey results, as well as to submit other work to our corpus (see https://imldresden.github.io/huis).
Figure 4: We present a taxonomy of key characteristics for mixed reality hybrid user interfaces with eight dimensions: configuration, temporal, relationship, range, device dependency, space, interaction dynamics, and anchoring. Interaction components are highlighted in green for content on a 2D screen and blue for content in a mixed reality environment.
3.6 How to Use This Survey
With our survey, we aim to reveal common patterns and identify shared dimensions within MR-HUIs. Our work can be helpful in three ways: (1) By classifying past research, our taxonomy of key characteristics provides a general framework that can be used to better describe and understand inherent properties of MR-HUIs. With these dimensions, we can share insights between systems with overlapping dimensions and identify related works within our literature corpus that may not share a common terminology. (2) By identifying current trends, our survey shows emerging trends in device combinations, use cases, contribution types, and evaluation strategies. Readers can understand how MR-HUIs are used, informing the design and evaluation of upcoming MR-HUIs. Finally, (3) our survey can be used to inspire future systems, serving as a roadmap to design the next generation of HUIs.
Throughout our reporting (Sec. 4 and 5), we denote the amount of literature in our corpus for each characteristic within a colored box mapped on a gradient with the total paper count (n=83) as the maximum value. Additionally, we provide a curated set of exemplary papers from our corpus for each characteristic, which can be used as a starting point for further reading. Since systems can be attributed to multiple characteristics, dimensions, or other categories, the sum of each category does not necessarily match the total number of records within our corpus.
@@ 4 A Taxonomy of Mixed Reality Hybrid User Interfaces
Based on our survey, we establish a taxonomy of key characteristics for MR-HUIs (Sec. 4.1) that can characterize existing research and inform new research. In addition, we describe emergent trends and opportunities (Sec. 4.2) and highlight edge cases (Sec. 4.3) of our attributes.
@@ 4.1 Taxonomy Dimensions
We adopt six cross-device characteristics of Brudy et al. [16] (Dimensions 1–6) and introduce two additional dimensions to represent better the design space of MR-HUIs. To better reflect the nuances of MR-HUI, we extend and reframe each dimension and explain them accordingly. Although we describe these dimensions as discrete characteristics, they should be seen as continuous spectra, where systems, configurations, and interaction techniques can dynamically span multiple facets.
@@
The configuration dimension describes how content and control are distributed across interaction components for a single user. Although we based our configurations on the cross-device taxonomy (i.e., mirrored, logical distribution, spatial distribution, migratory interfaces), we identified six configurations unique to MR-HUIs, namely asymmetrical mirrored, symmetrical mirrored, remote control, dynamic lens, augmented displays, and VESADs.
@@
configurations duplicate content in different interaction components. Given the perceptual differences of components in MR-HUIs, we distinguish between symmetric mirror and asymmetric mirror configurations.
$\triangleright$ Symmetric Mirror 7
configurations duplicate content between interaction components, with each interaction component displaying the exact same view of information. This can reduce complexity when interacting with virtual content such as 3D sketching [3, 34] or immersive analytics [67, 18]. For example, a HUI can provide contextual connection based on users’ surroundings in MR while providing familiar input on the 2D interaction component.
$\triangleright$ Asymmetric Mirror 8
configurations show the same view of information but make use of the additional perceptual dimensions offered by MR interaction components: For example, this can be used to display a full 3D model or visualization in the MR environment, which is mirrored to an orthographic front view on the 2D interaction component [141, 73, 132].
Logical Distribution 54
describes configurations where content and control are distributed according to each interaction component’s strength, usually involving a mutually exclusive allocation of responsibilities between interaction components. Most systems in our corpus involve some kind of logical distribution, such as offloading text input [55, 158, 8]), shared and personal space [121, 76], or general application-control [124, 121, 103]) to 2D interaction components, while the MR environment is used to display content in-situ.
$\triangleright$ Remote Control 6
describes a stricter subset of logical distribution, where the 2D interaction component provides an alternative (but not exclusive) control over content in the MR environment. This can be useful for providing direct interactions when close by and indirect interaction from farther away (e.g., during 3D sketching [34]) or providing a more ergonomic option [41].
@@
describes configurations that deliberately spread content across different spatial locations in the continuous real-world space. Such a distribution can be achieved through different means, described as dynamic lenses, augmented displays, or VESADs. Here, the content is aligned to the 2D interaction component (i.e., augmented display, VESAD) or reacts to the position of the 2D interaction component (i.e., dynamic lens).
$\triangleright$ Dynamic Lens 6
allows 2D interaction components to act as a dynamic peephole [99] into a larger information space. While 2D interaction components provide a constrained view, the MR environment is unrestricted and can use the real environment. This requires the 2D interaction component to be spatially aware, enabling, for example, 3D slicing showing a cross-section of a 3D model on the 2D interaction component [96, 85, 141].
$\triangleright$ Augmented Displays 15
use the unrestricted visual output of an MR environment to extend a 2D interaction component beyond its potential visual and interaction capabilities, acting as a 3D augmentation that is attached to the 2D interaction component [87, 119, 27, 33] . This category was initially defined by Reipschläger et al. [118, 120] as “seamless combination of high resolution touch and pen enabled displays with head-coupled Augmented Reality”.
$\triangleright$ VESADs 16
(“Virtually Extended Screen-Aligned Displays”) were initially defined by Normand and McGuffin [104] as a virtual AR screen “that is centered on, and co-planar with, a smartphone”. The content is strictly aligned to a 2D interaction component as a seamless display extension [87, 70, 47], representing a subset of augmented displays. For example, VESADs can be useful for offloading menu elements into MR to save screen real estate on a 2D interaction component [104, 120, 15] or provide alternative interaction capabilities [15]. To better represent the diversity of real-world multi-display configurations, we also include configurations that are aligned to 2D interaction components but not necessarily co-planar, such as extending 2D interaction components with angled virtual 2D screens [117, 110, 109].
Migratory Interfaces 11
enable users to transfer their content or workflow from one device to another. Although originally considered asynchronous within the cross-device taxonomy [16], MR-HUIs can also utilize the MR environment to seamlessly transfer content between 2D interaction components [133] or between a 2D and 3D interaction component [148, 150, 2, 153]. In contrast, their asynchronous usage can take advantage of different environments to best support a holistic workflow (e.g., by switching between VR and desktop environment [66, 1, 24]) while still behaving as one unified, continuous system.
@@
Prior literature [16, 13, 69] classifies hybrid and cross-device systems as synchronous or asynchronous. Due to the diversity of interaction components within MR-HUIs, we adopted the suggestion by Bornik et al. [13] to further differentiate between parallel and serial usage of interaction components. Fully asynchronous usage of interaction components was further classified as exclusive for cases where the usage of one interaction component rules out the usage of another (e.g., due to spatial or time-related restrictions).
@@
Parallel 67 / 43
usage indicates that multiple interaction components are used simultaneously, for example, when interacting with one interaction component while observing the output on another interaction component [27], transferring content across devices [67], or extending a 2D interaction component in the MR environment [87, 119, 70].
Serial 36 / 6
usage indicates that multiple interaction components are used one immediately after another, such as selecting objects of interest in the MR environment and then editing them on a 2D interaction component [18]. Here, users have immediate access to all interaction components but can focus on only one interaction component at a time to reduce information overload or divide responsibilities between 2D and 3D interaction [34, 46, 13].
Exclusive 12 / 4
usage describes asynchronous systems where different interaction components cannot be used simultaneously as part of one workflow but have to be used in sequence. This can be useful to bridge the gap between traditional 2D computing environments and VR environments  [66, 1, 24, 130].

### v2

@@ 2.3 Characterizing Hybrid User Interfaces
Research on HUIs shares several common themes (e.g., complementarity, heterogeneity), yet the term remains technology-driven and potentially misunderstood: A large number of prior HUIs describe the term as a combination of 2D and 3D technology [bornik2006hybrid, elmqvist2011distributed, hinckley1994survey], including the initial system demonstration by Feiner and Shamash [feiner1991hybrid]. Moreover, the term HUI was previously also used for different constellations of IO components (e.g., desktop combined with tangibles [strawhacker2015want], mobile devices [bessghaier2017usability], or conversational interfaces [liu2022conversationbased]), indicating wider applicability. For the purpose of our survey, we modernize the definition of HUIs by Feiner and Shamash [feiner1991hybrid] within the current research landscape:
@@
Hybrid User Interfaces are an area of cross-device computing that leverages distinct benefits of heterogeneous components with input and output (IO components).
@@
In contrast to its initial definition, we explicitly concentrate on conceptual device capabilities instead of specific technologies to avoid being limited by current hardware capabilities, but are intentionally vague about possible device combinations (e.g., see cross-device interaction [brudy2019crossdevice] for a potential ontology). To this end, we use the term IO component throughout this work to refer to a standalone device set with the input and output capabilities necessary for interacting with a given application (e.g., desktop with mouse and keyboard, VR HWD with controllers).
@@
While this definition highlights the broad potential design space of HUIs, it also lacks specificity and, therefore, limits its utility for our survey: For example, a combination of tangibles and a desktop computer might be considered a HUI and likewise, a combination of an AR HWD with a tablet—increasing the blurriness of the research stream instead of increasing its focus. Therefore, to explore this vast design space and gain concrete insights into commonalities of HUIs, we focus our work on the specific HUI subset combining 2D with MR-enabled IO components. This combination captures our interest, as it was initially presented by Feiner and Shamash [feiner1991hybrid], is most prevalent in prior HUI literature [satkowski2023fantastic], and also reflects on work presented at the IEEE ISMAR workshop on HUIs (2023) [hubenschmid2023hybrid], constituting the most recent and comprehensive outlet for work on HUIs. We thereby explicitly exclude the vast research area of tangible interaction [ishi2003tangible] or nascent research areas (e.g., brain–computer interfaces [chun2016bci], Internet of Things [sanctorum2019unifying]) from this subset, thereby increasing the specificity of our survey.
@@ 2.4 Attributes of Mixed Reality Hybrid User Interfaces
We framed our work on the contemporary and most prevalent subset of HUIs: a combination of 2D IO components with MR environments. We derive three attributes based on previous usage of the term to further guide our literature review. Please note that the first two attributes are indicative of the broader area of HUIs, while the last attribute specifies the constraints of MR-HUIs.
Multiple IO components in heterogeneous roles.  HUIs typically combine multiple IO components where each fulfills a need that is not adequately addressed by other components. Although each IO component is self-contained, a HUI is deliberately spread across multiple components to intentionally “take advantage of the strong points of each” [feiner1991hybrid]. We also see the potential beyond technological aspects and broaden our scope to include combinations of heterogeneous roles.
IO components are codependent.  A HUI is composed of multiple IO components, but the real power does not emerge from any individual IO component, but from the interaction of all of them [weiser1999computer]. A HUI thus acts as one holistic application from the user’s perspective. The deliberate spread of responsibilities across IO components requires a certain degree of co-dependency between IO components (i.e., a system may become nonfunctional without all IO components present). We examine both synchronous (i.e., using IO components in parallel) and asynchronous usage (i.e., using IO components in sequence) [bornik2006hybrid, hubenschmid2021asynchronous].
Complementing 2D with MR IO components.  A common aspect found in prior HUI literature is to “pair 3D perception and direct 3D interaction with 2D system control and precise 2D interaction” [bornik2006hybrid]. Given the (at times) conflicting needs of 2D and 3D spaces, combining 2D and MR IO components can yield a superior result. For the remainder of this work, we refer to such IO components as either 2D IO component or MR IO component.
@@ 3 Review Methodology
We look at research that matches our previously specified attributes to obtain corpus of MR-HUI publications. This section describes the process for identifying, filtering, and analyzing relevant publications, following the PRISMA [page2021prisma] guidelines. We present the general search strategy (Sec. 3.1), selection process (Sec. 3.2), and data extraction (Sec. 3.3). Moreover, the limitations of our procedure (Sec. 3.4) and possible extensions of the survey (Sec. 3.5) are highlighted. Finally, we describe how this survey can be used by other researchers and practitioners (Sec. 3.6). The complete survey corpus, its coding, and the scripts used to prepare and analyze the data can be found in the supplementary material.
Figure 2: Overview of our reviewing process including paper counts, following PRISMA. Records were rejected once a single exclusion criterium was fulfilled, yet, some records potentially fulfilled multiple criteria (e.g., survey papers [auda2023scoping]). The initial corpus of papers (n=1235) including all codes is part of the supplemental material of this submission.
@@ 3.1 Search Strategy
Identification via databases.  Our survey focuses on publications that combine 2D IO components with MR environments. As the term “HUI” is not consistently used in previous work, we used adjacent terms and their synonyms (see Sec. 2.2) to build a query that captures a broader range of potential publications that present HUI applications.
@@
The query was constructed using the expression: Set A or (Set B and Set C). We focused on archival, peer-reviewed publications, including full and short papers in journals and conferences, as well as book chapters, workshop submissions, posters, and works in progress. We searched in common digital libraries for publications in Human–Computer Interaction and Visualization, namely IEEE Xplore, ACM Digital Library, Springer Link, and Wiley Online Library.The searches with the described query [^4] The specific search syntax for each library can be found in the supplemental material. in these libraries [^5] The advanced search of the Springer library only allows searching in the title and keywords and not in the abstract. resulted in 438 publications (see Fig. 2). The cut-off date for all searches was December 11th, 2025.
Identification via other methods.  In addition to the query search—and in line with PRISMA guidelines—we selected papers through three other methods (see Fig. 2). First, we selected all papers referring to seminal works of adjacent research areas (i.e., [elmqvist2011distributed, zagermann2022complementary, brudy2019crossdevice, auda2023scoping, mayer2024crossing, reipschlager2020augmented]) or to the original definition of HUIs [feiner1991hybrid], through a Google Scholar search using “Publish or Perish” [harzing2007publish]. Second, we manually added further publications that were part of the authors’ paper collection, discussed among four authors, and eventually identified as potentially relevant. Some of these handpicked papers did not appear through other search strategies as their individual focus was not the technological setting but the investigated use case (e.g., studying window management of virtual displays [pavanatto2024multiple]). Finally, we selected all publications presented at the IEEE ISMAR 2023 Workshop on Hybrid User Interfaces [hubenschmid2023hybrid], as contributions to this workshop have explicitly been peer-reviewed in the context of HUIs. The three methods resulted in 797 publications (see Fig. 2). The cut-off date was December 11th, 2025.
@@ 3.2 Selection Process
After retrieving the initial corpus of papers (n = 1235), we preprocessed the different output formats and merged them into one table. Three authors screened the reviewed corpus and filtered out duplicates and papers unrelated to MR. In addition, workshop proposals, complete books, dissertations, or conference proceedings were also removed.
Subsequently, three authors checked the eligibility of the remaining publications (n = 503). For that, the following exclusion criteria (based on Sec. 2.4) were defined:
  • The publication is a duplicate (content-wise) of another (e.g., demonstration [reichherzer2021secondsighta] of system [reichherzer2021secondsight]).
  • No prototype or system was presented (e.g., discussion of opportunities and challenges [zaky2023opportunities, zagermann2023challenges, krug2023discussing, reinschluessel2023exploring]).
  • Only a technological basis is described (e.g., frameworks [hubenschmid2023colibri, sanctorum2019unifying]).
  • The publication only presents a summary of systems already in our corpus (e.g., position papers [feiner2000environment, feiner1999importance]).
  • A pure MR system without other IO components is described (e.g., combining AR and VR environments [cho2023time, bambusek2023how]).
  • Other device capabilities are only used as input or output modality (i.e., IO components are not standalone, e.g., attaching input sensors to an AR HWD [manuri2015preliminary, das2023fingerbutton]).
  • The publication only presents a collaboration across asymmetric devices (e.g., one user on a 2D IO component, another in an MR environment [olin2021designing, lunding2022exposar]).
During both selection steps, we split the data set between authors and discussed entries if the decision was unclear. This resulted in 116 relevant publications within our corpus.
3.3 Data Extraction
Since we see HUIs as a sub-category of cross-device interaction, we started with aspects defined by Brudy et al. [brudy2019crossdevice] in their survey. This was extended with common HCI metadata, such as the contribution type [wobbrock2016research] or the evaluation strategy [ledo2018evaluation]. Furthermore, we recorded each publication’s challenges, future work, use case, and devices and terminology used. To validate the initial set of codes, we randomly selected 10 papers from our corpus that were coded by three authors. This allowed us to (1) verify if the categorization works, (2) add missing codes, and (3) establish a common ground between authors. With the categories finalized, four authors coded the remaining papers. To ensure the rigor and reliability of the analysis, we employed researcher triangulation: two authors independently coded all publications, and each paper in the final corpus was reviewed by at least two researchers. No author assessed the relevance of their own work. After each paper was coded, the same four authors discussed each publication and clarified conflicts in a coding meeting. They further clustered challenges, use cases, and terminology used throughout the corpus.
@@ 3.4 Limitations and Further Considerations
Strict Eligibility Criteria.  We searched for a specific type of device combination (i.e., MR-enabling devices and conventional 2D displays), leading to strict eligibility criteria (see Sec. 2.4). To avoid being overly restrictive, we decided to integrate papers that meet our characteristics but leave room for interpretation, especially system papers representing potential directions for future research. We discuss these edge cases and their implications in Sec. 4.3. We acknowledge that our corpus is not exhaustive with regard to possible HUI systems caused by the use of said criteria and our intentional focus on MR-HUI. Yet, we see our corpus as a representative and substantial set of research on MR-HUI, showcasing trends that can be generalized to the larger set of papers not captured by our query. Although no qualitative analysis can guarantee a single definitive taxonomy, we applied an overall methodology that strengthens the validity of our results and ensures that the resulting structure is both rigorous and comprehensive.
@@ 3.5 Dissemination and Extension
We have made our literature corpus available using the Indy Survey Tool [crnovrsanin2023indy] as a GitHub project that hosts an interactive website for users to explore and filter our survey results, as well as to submit other work to our corpus (see https://imldresden.github.io/huis).
Figure 4: We present a taxonomy of key characteristics for mixed reality hybrid user interfaces with eight dimensions: configuration, temporal, relationship, range, device dependency, space, interaction dynamics, and anchoring. IO components are highlighted in green for content on a 2D screen and blue for content in a mixed reality environment.
3.6 How to Use The Following Survey
Our survey aims to reveal common patterns and identify shared dimensions within MR-HUIs. Our work can be helpful in three ways: (1) By classifying past research, our following taxonomy of key characteristics provides a general framework that can be used to describe and understand inherent properties of MR-HUIs. (2) By identifying current trends, our survey shows emerging trends in device combinations, use cases, contribution types, and evaluation strategies. Readers can understand how MR-HUIs are used, informing the design and evaluation of upcoming MR-HUIs. Finally, (3) our survey can inspire future systems, serving as a roadmap for the next generation of HUIs.
Throughout our reporting (Sec. 4 and 5), we denote the amount of literature in our corpus for each characteristic within a colored box mapped on a gradient with the total paper count (n=116) as the maximum value. Additionally, we provide a curated set of exemplary papers from our corpus for each characteristic, which can be used as a starting point for further reading. Since systems can be attributed to multiple characteristics, dimensions, or other categories, the sum of each category does not necessarily match the total number of records within our corpus.
@@ 4 A Taxonomy of Mixed Reality Hybrid User Interfaces
Based on our survey, we establish a taxonomy of key characteristics for MR-HUIs (Sec. 4.1) that can characterize existing research and inform new research. In addition, we describe emergent trends and opportunities (Sec. 4.2) and highlight edge cases (Sec. 4.3) of our attributes characterizing MR-HUIs.
@@ 4.1 Taxonomy Dimensions
We adopt six cross-device characteristics of Brudy et al. [brudy2019crossdevice] (Dimensions 1–6) and introduce two additional dimensions to better represent the design space of MR-HUIs. To reflect the nuances of MR-HUI, we extend and reframe each dimension and explain them accordingly. Although we describe these dimensions as discrete characteristics, they should be seen as continuous spectra, where systems, configurations, and interaction techniques can dynamically span multiple facets.
@@
The configuration dimension describes how content and control are distributed across IO components for a single user. Although we based our configurations on the cross-device taxonomy (i.e., mirrored, logical distribution, spatial distribution, migratory interfaces), we identified six configurations unique to MR-HUIs, namely asymmetrical mirrored, symmetrical mirrored, remote control, dynamic lens, augmented displays, and VESADs.
@@
configurations duplicate content in different IO components. Given the perceptual differences of components in MR-HUIs, we distinguish between symmetric mirror and asymmetric mirror configurations.
$\triangleright$ Symmetric Mirror 14
configurations duplicate content between IO components, with each IO component displaying the exact same view of information. This can reduce complexity when interacting with virtual content such as 3D sketching [arora2018symbiosissketch, drey2020vrsketchin] or immersive analytics [hubenschmid2021stream, butscher2018clusters]. For example, a HUI can provide contextual connection based on users’ surroundings in MR while providing familiar input on the 2D IO component.
$\triangleright$ Asymmetric Mirror 15
configurations show the same view of information but make use of the additional perceptual dimensions offered by MR IO components: For example, this can be used to display a full 3D model or visualization in the MR environment, which is mirrored to an orthographic front view on the 2D IO component [szalavari1997personal, iquiapaza2024dear, sereno2022hybrid].
Logical Distribution 67
describes configurations where content and control are distributed according to each IO component’s strength, usually involving a mutually exclusive allocation of responsibilities between IO components. Most systems in our corpus involve some kind of logical distribution, such as offloading text input [grubert2023text, zhu2020bishare, berns2019myr]), shared and personal space [rekimoto1999augmented, james2023evaluating], or general application-control [satkowski2021insitu, rekimoto1999augmented, nebeling2020mrat]) to 2D IO components, while the MR environment is used to display content in-situ.
$\triangleright$ Remote Control 12
describes a stricter subset of logical distribution, where the 2D IO component provides an alternative (but not exclusive) control over content in the MR environment. This can be useful for providing direct interactions when close by and indirect interaction from farther away (e.g., during 3D sketching [drey2020vrsketchin]) or providing a more ergonomic option [estalagem2024wearable].
@@
describes configurations that deliberately spread content across different spatial locations in the continuous real-world space. Such a distribution can be achieved through different means, described as dynamic lenses, augmented displays, or VESADs. Here, the content is aligned to the 2D IO component (i.e., augmented display, VESAD) or reacts to the position of the 2D IO component (i.e., dynamic lens).
$\triangleright$ Dynamic Lens 10
allows 2D IO components to act as a dynamic peephole [mehra2006navigating] into a larger information space. While 2D IO components provide a constrained view, the MR environment is unrestricted and can use the real environment. This requires the 2D IO component to be spatially aware, enabling, for example, 3D slicing showing a cross-section of a 3D model on the 2D IO component [luo2021exploring, krug2022clear, szalavari1997personal].
$\triangleright$ Augmented Displays 18
use the unrestricted visual output of an MR environment to extend a 2D IO component beyond its potential visual and interaction capabilities, acting as a 3D augmentation that is attached to the 2D IO component [langner2021marvis, reipschlager2021personal, chulpongsatorn2023holotouch, dedual2011creating] . This category was initially defined by Reipschläger et al. [reipschlager2020augmented, reipschlager2018debugar] as “seamless combination of high resolution touch and pen enabled displays with head-coupled Augmented Reality”.
$\triangleright$ VESADs 19
(“Virtually Extended Screen-Aligned Displays”) were initially defined by Normand and McGuffin [normand2018enlarging] as a virtual AR screen “that is centered on, and co-planar with, a smartphone”. The content is strictly aligned to a 2D IO component as a seamless display extension [langner2021marvis, hubenschmid2023smartphone, feiner1991hybrid], representing a subset of augmented displays. For example, VESADs can be useful for offloading menu elements into MR to save screen real estate on a 2D IO component [normand2018enlarging, reipschlager2018debugar, brasier2021arenhanced] or provide alternative interaction capabilities [brasier2021arenhanced]. To better represent the diversity of real-world multi-display configurations, we also include configurations that are aligned to 2D IO components but not necessarily co-planar, such as extending 2D IO components with angled virtual 2D screens [reipschlager2019designar, pavanatto2021we, pavanatto2024multiple].
Migratory Interfaces 22
enable users to transfer their content or workflow from one device to another. Although originally considered asynchronous within the cross-device taxonomy [brudy2019crossdevice], MR-HUIs can also utilize the MR environment to seamlessly transfer content between 2D IO components [serrano2015gluey] or between a 2D and 3D IO component [wang2024user, wentzel2024switchspace, aigner2023cardiac, wu2020megereality]. In contrast, their asynchronous usage can take advantage of different environments to best support a holistic workflow (e.g., by switching between VR and desktop environment [hubenschmid2022relive, chan2023single]) while still behaving as one unified, continuous system.
@@
Prior literature [brudy2019crossdevice, bornik2006hybrid, hubenschmid2021asynchronous] classifies hybrid and cross-device systems as synchronous or asynchronous. Due to the diversity of IO components within MR-HUIs, we adopted the suggestion by Bornik et al. [bornik2006hybrid] to further differentiate between parallel and serial usage of IO components. Fully asynchronous usage of IO components was further classified as exclusive for cases where the usage of one IO component rules out the usage of another (e.g., due to spatial or time-related restrictions).
@@
Parallel 94 / 65
usage indicates that multiple IO components are used simultaneously, for example, when interacting with one IO component while observing the output on another IO component [chulpongsatorn2023holotouch], transferring content across devices [hubenschmid2021stream], or extending a 2D IO component in the MR environment [langner2021marvis, reipschlager2021personal, hubenschmid2023smartphone].
Serial 45 / 9
usage indicates that multiple IO components are used one immediately after another, such as selecting objects of interest in the MR environment and then editing them on a 2D IO component [butscher2018clusters]. Here, users have immediate access to all IO components but can focus on only one IO component at a time to reduce information overload or divide responsibilities between 2D and 3D interaction [drey2020vrsketchin, feiner1997touring, bornik2006hybrid].
Exclusive 17 / 5
usage describes asynchronous systems where different IO components cannot be used simultaneously as part of one workflow but have to be used in sequence. This can be useful to bridge the gap between traditional 2D computing environments and VR environments  [hubenschmid2022relive, chan2023single, schroder2023collaborating].